# Optimal Conformal Prediction, E-values, Fuzzy Prediction Sets and Subsequent Decisions

### v1

@@ 2.2 The non-coverage event as a test
A prediction confidence set $C_{\alpha}$ is equivalent to the test $\phi:\mathcal{Z}^{n+1}\to\{0,1\}$, defined as $\phi(z^{n+1})=\mathbb{I}\{z_{n+1}\in C_{\alpha}^{z^{n}}\}$. The confidence set is valid at level $\alpha$ for $\mathcal{P}$ if and only if the test $\phi$ is.
Proof.
Given a prediction confidence set $C_{\alpha}$, we define the test $\phi$ so that it rejects when $C_{\alpha}^{z^{n}}$ does not cover $z_{n+1}$:
$\phi(z^{n+1})=\mathbb{I}\{z_{n+1}\not\in C_{\alpha}^{z^{n}}\}.$ (5)
Conversely, given the test $\phi$ we may recover the confidence set $C_{\alpha}$ by evaluating it at different points $(z_{1},\dots,z_{n},z_{n+1})$ in the sample space $\mathcal{Z}^{n+1}$:
$C_{\alpha}^{z^{n}}=\{z\in\mathcal{Z}:\phi(z_{1},\dots,z_{n},z)=0\}.$
As we may convert from one to the other without loss, the test $\phi$ and confidence set $C_{\alpha}$ are equivalent.
The equivalence of their validity directly follows from rewriting the coverage guarantee as
$P(Z_{n+1}\not\in C_{\alpha}^{Z^{n}})\leq\alpha,\textnormal{ for every }P\in\mathcal{P},$
[…]
2.3 From confidence width to power
By Theorem 1, constructing a valid prediction confidence set for $\mathcal{P}$ is equivalent to constructing a valid test for $\mathcal{P}$. It then remains to select a test that yields a ‘good’ or even optimal confidence set in some appropriate sense. Here, one may hope that an optimal confidence set corresponds to a test that is most powerful against some alternative. We show this is indeed the case, by deriving the alternative hypothesis under which the test should be most powerful, in order for our confidence set to be ‘as small as possible’.
To formalize what we mean with ‘as small as possible’, let $\mu$ denote some unsigned measure on $\mathcal{Z}$ selected to measure the size of our confidence set. For the sake of exposition, let us assume we have access to the true distribution $P_{*}^{Z^{n}}$ of $Z^{n}$. It then seems natural to formalize our problem as minimizing the expected $\mu$-size of our confidence set under $P_{*}^{Z^{n}}$:
@@ 2.3 From confidence width to power
over valid confidence sets $C_{\alpha}$. For example, if $\mathcal{Z}$ is finite then a natural choice for $\mu$ may be the counting measure, which leads to the minimization of the number of elements in our confidence set:
@@
In Theorem 2, we show that $C_{\alpha}$ minimizes the expected size of the confidence set if and only if its corresponding test $\phi$ is most powerful against the alternative $Q=P_{*}^{Z^{n}}\times\mu$. In the result, we even show that the choice of $\mu$ is allowed to depend on $Z^{n}$.
@@
if and only if its corresponding test $\phi$ is most powerful against the alternative $Q=P_{*}^{Z^{n}}\otimes\mu^{|Z^{n}}$.
@@
This result follows immediately from the equivalence between the confidence set $C_{\alpha}$ and its test $\phi$:
@@
$=\operatorname*{arg\,min}_{\phi}\int_{\mathcal{Z}^{n}}\int_{\mathcal{Z}}\mathbb{I}\{Z_{n+1}\in C_{\alpha}^{Z^{n}}\}\,d\mu^{|Z^{n}}\,dP_{*}^{Z^{n}}$
$=\operatorname*{arg\,min}_{\phi}\int_{\mathcal{Z}^{n+1}}(1-\phi)\,d(P_{*}^{Z^{n}}\otimes\mu^{|Z^{n}})$
$=\operatorname*{arg\,max}_{\phi}\int_{\mathcal{Z}^{n+1}}\phi\,d(P_{*}^{Z^{n}}\otimes\mu^{|Z^{n}})$
$=\operatorname*{arg\,max}_{\phi}\mathrm{E}^{Q}[\phi],$
@@
In practice, the true distribution $P_{*}^{Z^{n}}$ is generally unknown, so that the optimal procedure under the true distribution cannot be used in practice. However, we must stress that the lack of such oracle knowledge does not endanger the validity of the procedure; it only harms its performance. Indeed, if we plug-in some guess for $P_{*}^{Z^{n}}$, then we would still minimize the size of the confidence set in expectation over our guess. As a consequence, we view the choice of plug-in for $P_{*}^{Z^{n}}$ as expressing the distribution of $Z^{n}$ under which we wish the confidence set to be small in expectation.
The fact that we have no oracle knowledge of $P_{*}^{Z^{n}}$ may be viewed as special case of a general problem in hypothesis testing: we can rarely hope to have a test that is optimal uniformly against every distribution in a composite alternative hypothesis. While there is no general-purpose tool to truly ‘solve’ this problem, it is well-studied and solutions exist for special cases. For example, in Section 2.4 we may get around this issue using the Hunt-Stein theorem. In Section 4, we use other techniques to reduce the hypothesis by using the structure of the problem. In case there is no structure available, a general-purpose tool is to construct a valid test against several (or all) possible guesses of $P_{*}^{Z^{n}}$, and choose $\phi$ as some desired probability mixture over these tests.
@@ 2.4 Examples: Gaussian location
In this section, we illustrate our methodology in Gaussian models. These examples may be viewed as an illustration, to show that our framework naturally yields familiar prediction confidence sets. In Example 1, $\mathcal{P}$ contains a single distribution and the observations are independent, so that there is nothing to learn from the $n$ observations; we indeed recover a classical confidence set. In Example 2, we find that our optimal confidence set coincides with the familiar prediction set for the Gaussian location family, which is normally derived through finding a pivot and imposing symmetry of the confidence set. In Example 3, we showcase the use of a kernel $\mu^{\mid Z^{n}}$.
@@
In case $\mathcal{P}$ contains just a single distribution, $\mathcal{P}=\{P\}$, then under independence of $Z^{n}$ and $Z_{n+1}$ there is nothing to learn from the first $n$ observations. As a consequence, the optimal (fuzzy) confidence set $C_{\alpha}$ does not depend on the first $n$ observations $z^{n}\in\mathcal{Z}^{n}$ and so is deterministic.
To illustrate this, let $\mathcal{Z}=\mathbb{R}$, and suppose our class of distributions contains a single multivariate Gaussian distribution $\mathcal{P}=\{\mathcal{N}(\mu 1_{n+1},\sigma^{2}I_{n+1})\}$, for known $\mu\in\mathbb{R}$ and $\sigma>0$. A natural goal would be to minimize the Lebesgue measure $\lambda(C_{\alpha}^{Z^{n}})$ of our confidence set in expectation. This would come down to choosing the alternative $Q=\mathcal{N}(\mu 1_{n},\sigma^{2}I_{n})\times\lambda$. Unfortunately, such a non-finite measure causes the optimization problem to be poorly defined. Luckily, this may be resolved by replacing it with the proper alternative $Q=\mathcal{N}(\mu 1_{n},\sigma^{2}I_{n})\times\mathcal{N}(\mu,\tau^{2})$, for $\tau>\sigma$. We choose $\tau>\sigma$ so that the likelihood ratio $\frac{d\mathcal{N}(\mu,\tau^{2})}{d\mathcal{N}(\mu,\sigma^{2})}$ is a monotone transformation of the one with the Lebesgue measure $\frac{d\lambda}{d\mathcal{N}(\mu,\sigma^{2})}$.
For the classical Neyman–Pearson objective $\mathrm{E}^{Q}[\phi]$, it is then straightforward to show that the optimal test is the level $\alpha$ Neyman–Pearson optimal test:
$\phi(z^{n+1})$ $=\mathbb{I}\left\{\frac{d\mathcal{N}(\mu,\tau^{2})}{d\mathcal{N}(\mu,\sigma^{2})}(z_{n+1})>\frac{d\mathcal{N}(\mu,\tau^{2})}{d\mathcal{N}(\mu,\sigma^{2})}(c_{1-\alpha/2})\right\}$
@@
Note that this test is independent of $z^{n}$. Hence, we may simply evaluate it at every $z_{n+1}\in\mathcal{Z}$, for arbitrary $z^{n}$, to invert the test in $\mathcal{Z}$ and obtain the confidence set
@@
which is simply the standard two-sided Gaussian confidence interval. Here, we indeed have that the probability $Z_{n+1}$ is in the confidence set is $1-\alpha$.
@@
In case $\mathcal{P}$ is composite, the first $n$ observations are helpful in narrowing down the true distribution $P^{*}\in\mathcal{P}$. As a consequence, the optimal confidence set $C_{\alpha}^{Z^{n}}$ will now depend on the first observations $Z^{n}$.
To illustrate this, let $\mathcal{Z}=\mathbb{R}$, and suppose our class of distributions contains i.i.d Gaussians with known variance $\sigma^{2}>0$, $\mathcal{P}=\{\mathcal{N}(\mu 1_{n+1},\sigma^{2}I_{n+1}):\mu\in\mathbb{R}\}$. We again wish to minimize the Lebesgue measure $\lambda(C_{\alpha}^{Z^{n}})$ of our confidence set in expectation over the true distribution. For this purpose, we pick the proper composite alternative $\mathcal{Q}=\{\mathcal{N}(\mu 1_{n},\sigma^{2}I_{n})\times\mathcal{N}(\mu,\tau^{2}):\mu\in\mathbb{R}\}$, where $\tau>\sigma$. Moreover, define $A=Z^{n}-\bar{Z}_{n}$ and $B=Z_{n+1}-\bar{Z}_{n}$, where $\bar{Z}_{n}$ is the sample mean of $Z^{n}$.
@@
Now $A$ has the same distribution for any $\mu$ under both $\mathcal{P}$ and $\mathcal{Q}$, which drops out in the likelihood ratio. Moreover, $B\sim\mathcal{N}(0,\sigma^{2}+\sigma^{2}/n)$ under $\mathcal{P}$ and $B\sim\mathcal{N}(0,\tau^{2}+\tau^{2}/n)$ under $\mathcal{Q}$. This reduces the likelihood ratio to
@@
For the classical Neyman–Pearson power-objective $\mathrm{E}^{Q}[\phi]$, the optimal e-value is the level $\alpha$ Neyman–Pearson optimal test
$\phi(Z^{n+1})$ $=\mathbb{I}\left\{\textnormal{LR}^{B}(B)>\textnormal{LR}^{B}(c_{1-\alpha/2})\right\},$
@@
Comparing this to (6), we see that the known mean $\mu$ is replaced by an estimator $\bar{Z}_{n}$, and that we have an additional $\sigma^{2}/n$ pop-up in the variance which is due to the estimation of $\bar{Z}_{n}$.
@@
Now suppose $\mathcal{P}$ contains a single distribution, $\mathcal{P}=\{\mathcal{N}(\mu 1_{n+1},\Sigma)\}$, where $\Sigma_{i,j}=\rho^{|i-j|}$, for all pairs $(i,j)$. This corresponds to an AR(1) model with autoregressive parameter $\rho$.
@@
$\phi(z^{n+1})$ $=\mathbb{I}\left\{\frac{d\mathcal{N}(\mu_{n},\tau^{2})}{d\mathcal{N}(\mu_{n},1)}(z_{n+1})>\frac{d\mathcal{N}(\mu_{n},\tau^{2})}{d\mathcal{N}(\mu_{n},1)}(c_{1-\alpha/2})\right\}$
@@
which may again be interpreted as minimizing the Lebesgue measure of the confidence set. The resulting confidence set becomes
@@
which is simply the standard two-sided Gaussian confidence interval for an AR(1) model.
3 Fuzzy confidence sets
3.1 Fuzzy confidence sets
Fuzzy confidence sets are intended to move beyond the binary inclusion-exclusion $Z_{n+1}\not\in C_{\alpha}^{Z^{n}}$ offered by a classical confidence set $C_{\alpha}^{Z^{n}}$. Fuzzy confidence set were originally proposed by Geyer and Meeden (2005) for confidence sets on parameter spaces, to help in expressing confidence in case of discrete data. Our contribution here is to show that they are of general interest, beyond solving discreteness issues.
To describe a fuzzy confidence set, it is helpful to reinterpret the confidence set $C_{\alpha}^{z^{n}}$ as a map

### v2

@@ 3.2 Prediction as a testing problem
A prediction set $C_{\alpha}^{Z^{n}}$ is equivalent to the test $C_{\alpha}$,
$C_{\alpha}(z^{n+1}):=\mathbb{I}\{z^{n+1}\not\in C_{\alpha}^{z^{n}}\}.$
The prediction set $C_{\alpha}^{Z^{n}}$ is valid at level $\alpha$ for the model $\mathcal{P}$ if and only if this test $C_{\alpha}$ is valid at level $\alpha$ for the hypothesis $\mathcal{P}$.
Remark 2 (Connection to classical testing).
At first glance the connection to testing may seem superficial, but we argue it is not. Recall that in classical hypothesis testing we hypothesize a model $\mathcal{P}$ (often called “the hypothesis”) and subsequently observe data. As we observe the data, we must reject the hypothesized model if the data and model are incompatible. In the construction of a prediction sets the roles are reversed: we fix a model $\mathcal{P}$ as the ground truth and hypothesize data $Z^{n+1}=z^{n+1}$. If the two are incompatible, we must now reject the hypothesized data. Extending this argument: if a part $Z^{n}=z^{n}$ of $Z^{n+1}$ is also observed alongside the model $\mathcal{P}$, we must reject the hypothesized part: $Z_{n+1}=z_{n+1}$.
3.3 Optimal prediction sets and power
By Theorem 1, constructing a prediction set is equivalent to constructing a test. It then remains to select a test that yields a ‘good’ or even optimal prediction set in some appropriate sense. One may hope that an optimal prediction set corresponds to a test that is most powerful against some alternative. We show this is indeed the case, by deriving the alternative hypothesis under which the test should be most powerful in order for our prediction set to be ‘as small as possible’.
To formalize what we mean with ‘as small as possible’, let $\mu$ denote some unsigned measure on $\mathcal{Z}$, selected to measure the size of $C_{\alpha}^{Z^{n}}$. For the sake of exposition, let us temporarily assume we have access to the true distribution $P_{*}^{Z^{n}}$ of $Z^{n}$. It then seems natural to formalize our problem as minimizing the expected $\mu$-size of $C_{\alpha}^{Z^{n}}$ under $P_{*}^{Z^{n}}$:
@@ 3.3 Optimal prediction sets and power
over valid prediction sets $C_{\alpha}$. For example, if $\mathcal{Z}$ is finite then a natural choice for $\mu$ may be the counting measure, which leads to the minimization of the number of elements in our prediction set:
@@
In Theorem 2, we show that $C_{\alpha}$ minimizes the expected size of $C_{\alpha}^{Z^{n}}$ if and only if its test-representation is most powerful against the alternative $Q=P_{*}^{Z^{n}}\otimes\mu$. In the result, we even show that the choice of $\mu$ may depend on $Z^{n}$.
@@
if and only if the test $C_{\alpha}$ is most powerful against the alternative $Q=P_{*}^{Z^{n}}\otimes\mu^{|Z^{n}}$.
@@
This result follows immediately from the equivalence between the prediction set $C_{\alpha}$ and its test representation $C_{\alpha}$:
@@
$=\operatorname*{arg\,min}_{C_{\alpha}}\int_{\mathcal{Z}^{n}}\int_{\mathcal{Z}}\mathbb{I}\{Z_{n+1}\in C_{\alpha}^{Z^{n}}\}\,d\mu^{|Z^{n}}\,dP_{*}^{Z^{n}}$
$=\operatorname*{arg\,min}_{C_{\alpha}}\int_{\mathcal{Z}^{n+1}}(1-C_{\alpha})\,d(P_{*}^{Z^{n}}\otimes\mu^{|Z^{n}})$
$=\operatorname*{arg\,max}_{C_{\alpha}}\int_{\mathcal{Z}^{n+1}}C_{\alpha}\,d(P_{*}^{Z^{n}}\otimes\mu^{|Z^{n}})$
$=\operatorname*{arg\,max}_{C_{\alpha}}\mathrm{E}^{Q}[C_{\alpha}],$
@@
3.4 Handling unknown $P_{*}^{Z^{n}}$
In practice, the true distribution $P_{*}^{Z^{n}}$ is generally unknown. Fortunately, we can just plug-in an educated guess or estimator $\widehat{P}^{Z^{n}}$ for $P_{*}^{Z^{n}}$, based on a separate dataset. The resulting prediction set will still be valid for $\mathcal{P}$. The only consequence is that it is optimal for the estimator $\widehat{P}^{Z^{n}}$, instead of for the true distribution $P_{*}^{Z^{n}}$. This mirrors the classical problem in hypothesis testing that if we do not ‘know’ the true alternative distribution in case of a composite alternative, then we can rarely hope to construct a test that is most powerful against it, except for extremely rare settings where there exist uniformly most powerful tests.
In case there is no structure available, a general-purpose tool is to construct a valid test against several (or all) possible guesses of $P_{*}^{Z^{n}}$, and choose our test as some (probability)-weighted average over these tests. However, such an average would usually take value in $[0,1]$, so that this is an example of a fuzzy test, which we treat in Section 4. Another strategy is to construct an optimal test against such a weighted average over different resulting choices of $Q$.
@@ 3.5 Examples: Gaussian location
In this section, we illustrate our methodology in Gaussian models. Here, we show that our framework naturally yields familiar prediction sets.
In Example 1, $\mathcal{P}$ contains a single distribution and the observations are independent, so that there is nothing to learn from the $n$ observations; we indeed recover a classical prediction set. In Example 2, we find that our optimal prediction set coincides with the familiar prediction set for the Gaussian location family, which is normally derived through finding a pivot and imposing symmetry of the prediction set. In Example 3, we showcase the use of a kernel $\mu^{\mid Z^{n}}$. In Example 4, we show what happens if we naively use the classical two-sided z-test to produce a prediction set, to highlight the importance to carefully select the alternative $Q$.
@@
In case $\mathcal{P}$ contains just a single distribution, $\mathcal{P}=\{P\}$, then under independence of $Z^{n}$ and $Z_{n+1}$ there is nothing to learn from the first $n$ observations. As a consequence, the optimal (fuzzy) prediction set $C_{\alpha}$ does not depend on the first $n$ observations $z^{n}\in\mathcal{Z}^{n}$ and so is deterministic.
To illustrate this, let $\mathcal{Z}=\mathbb{R}$, and suppose our class of distributions contains a single multivariate Gaussian distribution $\mathcal{P}=\{\mathcal{N}(\mu 1_{n+1},\sigma^{2}I_{n+1})\}$, for known $\mu\in\mathbb{R}$ and $\sigma>0$. A natural goal would be to minimize the Lebesgue measure $\lambda(C_{\alpha}^{Z^{n}})$ of our prediction set in expectation. This would come down to choosing the alternative $Q=\mathcal{N}(\mu 1_{n},\sigma^{2}I_{n})\times\lambda$. Unfortunately, such a non-finite measure causes the optimization problem to be poorly defined. Luckily, this may be resolved by replacing it with the proper alternative $Q=\mathcal{N}(\mu 1_{n},\sigma^{2}I_{n})\times\mathcal{N}(\mu,\tau^{2})$, for $\tau>\sigma$. We choose $\tau>\sigma$ so that the likelihood ratio $\frac{d\mathcal{N}(\mu,\tau^{2})}{d\mathcal{N}(\mu,\sigma^{2})}$ is a monotone transformation of the one with the Lebesgue measure $\frac{d\lambda}{d\mathcal{N}(\mu,\sigma^{2})}$.
For the classical Neyman–Pearson objective $\mathrm{E}^{Q}[C_{\alpha}]$, it is then straightforward to show that the optimal test is the level $\alpha$ Neyman–Pearson optimal test:
$C_{\alpha}(z^{n+1})$ $=\mathbb{I}\left\{\frac{d\mathcal{N}(\mu,\tau^{2})}{d\mathcal{N}(\mu,\sigma^{2})}(z_{n+1})>\frac{d\mathcal{N}(\mu,\tau^{2})}{d\mathcal{N}(\mu,\sigma^{2})}(c_{1-\alpha/2})\right\}$
@@
Note that this test is independent of $z^{n}$. Hence, we may simply evaluate it at every $z_{n+1}\in\mathcal{Z}$, for arbitrary $z^{n}$, to invert the test in $\mathcal{Z}$ and obtain the prediction set
@@
which is simply the standard two-sided Gaussian prediction interval. Here, we indeed have that the probability $Z_{n+1}$ is in the prediction set is $1-\alpha$.
@@
In case $\mathcal{P}$ is composite, the first $n$ observations are helpful in narrowing down the true distribution $P_{*}^{Z^{n+1}}\in\mathcal{P}$. As a consequence, the optimal prediction set $C_{\alpha}^{Z^{n}}$ will now depend on the first observations $Z^{n}$.
To illustrate this, let $\mathcal{Z}=\mathbb{R}$, and suppose our class of distributions contains i.i.d Gaussians with known variance $\sigma^{2}>0$, $\mathcal{P}=\{\mathcal{N}(\mu 1_{n+1},\sigma^{2}I_{n+1}):\mu\in\mathbb{R}\}$. We again wish to minimize the Lebesgue measure $\lambda(C_{\alpha}^{Z^{n}})$ of our prediction set in expectation over the true distribution. For this purpose, we pick the proper composite alternative $\mathcal{Q}=\{\mathcal{N}(\mu 1_{n},\sigma^{2}I_{n})\times\mathcal{N}(\mu,\tau^{2}):\mu\in\mathbb{R}\}$, where $\tau>\sigma$. Moreover, define $A=Z^{n}-\bar{Z}_{n}$ and $B=Z_{n+1}-\bar{Z}_{n}$, where $\bar{Z}_{n}$ is the sample mean of $Z^{n}$.
@@
Now, $A$ has the same distribution for any $\mu$ under both $\mathcal{P}$ and $\mathcal{Q}$, which drops out in the likelihood ratio. Moreover, $B\sim\mathcal{N}(0,\sigma^{2}+\sigma^{2}/n)$ under $\mathcal{P}$ and $B\sim\mathcal{N}(0,\tau^{2}+\tau^{2}/n)$ under $\mathcal{Q}$. This reduces the likelihood ratio to
@@
For the classical Neyman–Pearson power-objective $\mathrm{E}^{Q}[C_{\alpha}]$, the optimal e-value is the level $\alpha$ Neyman–Pearson optimal test
$C_{\alpha}(Z^{n+1})$ $=\mathbb{I}\left\{\textnormal{LR}^{B}(B)>\textnormal{LR}^{B}(c_{1-\alpha/2})\right\},$
@@
Comparing this to (6), we see that the known mean $\mu$ is replaced by an estimator $\bar{Z}_{n}$, and that we have an additional $\sigma^{2}/n$ appear in the variance which is due to the uncertainty introduced by the estimation of $\mu$ by $\bar{Z}_{n}$.
@@
Now suppose $\mathcal{P}$ contains a single distribution, $\mathcal{P}=\{\mathcal{N}(\mu 1_{n+1},\Sigma)\}$, where $\Sigma_{i,j}=\rho^{|i-j|}$, for all pairs $(i,j)$. This corresponds to an AR(1) model with autoregressive parameter $\rho\in(-1,1)$.
@@
$C_{\alpha}(z^{n+1})$ $=\mathbb{I}\left\{\frac{d\mathcal{N}(\mu_{n},\tau^{2})}{d\mathcal{N}(\mu_{n},1)}(z_{n+1})>\frac{d\mathcal{N}(\mu_{n},\tau^{2})}{d\mathcal{N}(\mu_{n},1)}(c_{1-\alpha/2})\right\}$
@@
which may again be interpreted as minimizing the Lebesgue measure of the prediction set. The resulting prediction set becomes
@@
which is simply the standard two-sided Gaussian prediction interval for an AR(1) model.
Example 4 (Prediction set based on two-sided z-test).
As highlighted in Theorem 1, any valid test produces a valid prediction set. The most popular test for Gaussian data is perhaps the two-sided z-test, so one may wonder what kind of prediction set it produces. Here, our model is $\mathcal{P}=\{\mathcal{N}(\mu 1_{n+1},\sigma^{2}I_{n+1})\}$, as this is the null hypothesis for which the two-sided z-test for the mean $\mu$ is valid.
Recall that the two-sided z-test rejects a hypothesis at level $\alpha$ if $\bar{Z}_{n+1}>\mu+\sigma c_{1-\alpha/2}/\sqrt{n+1}$, where $c_{1-\alpha/2}$ is the standard Gaussian’s $(1-\alpha/2)$-quantile. From this rejection event, we then recover the prediction set as the values of $Z_{n+1}=z_{n+1}$ for which the test does not reject:
$[\mu+n(\mu-\bar{Z}_{n})$- $\sqrt{n+1}\sigma c_{1-\alpha/2},$
$\mu+n(\mu-\bar{Z}_{n})+\sqrt{n+1}\sigma c_{1-\alpha/2}].$
As the model $\mathcal{P}$ is the same as the one we considered in Example 1, we may directly compare the prediction sets. In this comparison, we see that the prediction set here depends on $Z^{n}$ and is centered around $\mu+n(\mu-\bar{Z}_{n})$ instead of $\mu$. Moreover, to compensate for the peculiar center, its radius is the much wider $\sqrt{n+1}\sigma c_{1-\alpha/2}$ instead of $\sigma c_{1-\alpha/2}$.
To conclude, while this is technically a perfectly valid prediction set, it is perhaps not the prediction set that one would desire to use in most applications. This highlights the importance of selecting an appropriate $Q$ to express our desire about the prediction set.
4 Fuzzy prediction sets
4.1 From classical to fuzzy prediction sets
Fuzzy prediction sets are intended to move beyond the binary inclusion-exclusion $Z_{n+1}\not\in C_{\alpha}^{Z^{n}}$ offered by a classical prediction set $C_{\alpha}^{Z^{n}}$.
To define a fuzzy prediction set, it is helpful to again consider the test-representation of a prediction set